OpenStudy (anonymous):

Use the given circumference to find the surface area of the spherical object. A pincushion with C = 12 cm The surface area of the pincushion is _____ cm^2 (Round to the nearest whole number as needed)

OpenStudy (anonymous):

Surface area of a sphere is given A=4πr2 So...you have to find the radius from the given circumference gimme a sec

OpenStudy (tkhunny):

Please use Circumference to calculate the Radius. Circumference / ??? = Diameter Diameter / ?? = Radius

OpenStudy (anonymous):

A=4πr2 C = 12 cm = diameter x pi (divide 12 by pi) C = 3.8 (about) x pi (divide 3.8 by 2 to get radius, since d=2r) r - 1.9 A = 4 x pi x 1.9^2 (plug it all in and calculate) A = 45.36 45

OpenStudy (anonymous):

So 45 would be the surface area of the pincushion

OpenStudy (tkhunny):

Well, except for the rounding issues. The actual answer is 46. Too much intermediate rounding. C = 12 D = C/pi = 12/pi r = D/2 = D/2 = (12/pi)/2 = 6/pi SA = 4*pi*r^2 = 4*pi*(6/pi)^2 = 4*6^2 * pi/(pi^2) = 4*36/pi = 144/pi = 45.8366 ==> 46 with no intermediate rounding.

OpenStudy (tkhunny):

Whoops. I left off all the units. Sorry about that. Don't feel like rewriting it.

OpenStudy (anonymous):

alright well i appricate it yall, thank you.